# Fine-grained Image Retrieval via Dual-Vision Adaptation

### v1

@@ Abstract.
Fine-Grained Image Retrieval (FGIR) faces challenges in learning discriminative visual representations to retrieve images with similar fine-grained features. Current leading FGIR solutions typically follow two regimes: enforce pairwise similarity constraints in the semantic embedding space, or incorporate a localization sub-network to fine-tune the entire model. However, such two regimes tend to overfit the training data while forgetting the knowledge gained from large-scale pre-training, thus reducing their generalization ability. In this paper, we propose a Dual-Vision Adaptation (DVA) approach for FGIR, which guides the frozen pre-trained model to perform FGIR through collaborative sample and feature adaptation. Specifically, we design Object-Perceptual Adaptation, which modifies input samples to help the pre-trained model perceive critical objects and elements within objects that are helpful for category prediction. Meanwhile, we propose In-Context Adaptation, which introduces a small set of parameters for feature adaptation without modifying the pre-trained parameters. This makes the FGIR task using these adjusted features closer to the task solved during the pre-training. Additionally, to balance retrieval efficiency and performance, we propose Discrimination Perception Transfer to transfer the discriminative knowledge in the object-perceptual adaptation to the image encoder using the knowledge distillation mechanism. Extensive experiments show that DVA has fewer learnable parameters and performs well on three in-distribution and three out-of-distribution fine-grained datasets.
Fine-grained image retrieval, dual vision adaptation, knowledge distillation
copyright: none
1. Introduction
Unlike general image retrieval, fine-grained image retrieval (FGIR) attempts to retrieve images belonging to the same subcategory as the query image from a database within a broader meta category (i.e., birds, cars) (Wei et al., 2017). It has been extensively applied across various domains, including intelligent transportation (Ramdurai, 2025) and biodiversity monitoring (Vendrow et al., 2024). However, FGIR presents significant challenges due to two inherent difficulties: i) subtle visual distinctions between different categories, and ii) significant appearance variations within the same category. This task usually requires models to simultaneously localize discriminative regions and identify minute visual differences, creating a dual optimization dilemma. Therefore, the key to FGIR is learning discriminative and generalizable embeddings to identify visually similar objects accurately.
Conventional FGIR methods typically adopt two primary strategies: encoding-based (Movshovitz-Attias et al., 2017; Teh et al., 2020; Ermolov et al., 2022; Lim et al., 2022) and localization-based (Zheng et al., 2018; Wang et al., 2023b; Wei et al., 2017; Wang et al., 2022). Encoding-based methods primarily learn image-level features but struggle to suppress background and irrelevant information interference. In contrast, localization-based methods enhance feature extraction within the encoder to capture subtle differences among subcategories, often by focusing on distinctive object regions. Although both strategies capture discriminative embedding learning for fine-grained retrieval, their dependence on full-parameter fine-tuning introduces prohibitive computational costs and limits cross-task generalization, as shown in Fig. 1(a). Simultaneously, fine-tuning the entire model destroys the knowledge in the large-scale pre-trained parameters and may lead to suboptimal convergence, which ultimately limits retrieval performance (Wang et al., 2023b). This raises the question: can discriminative representations be learned for FGIR tasks without fine-tuning the model’s pre-trained parameters?
@@ 1. Introduction
Figure 1. (a) previous fine-tuning methods. (b) our dual-vision adaptation method. Our approach designs the collaborative sample and feature adaptation to exploit category-specific differences. This dual strategy enables model to sustain broad representation capabilities from pre-training data while dynamically adjusting its adaptability to fine-grained data.
@@
Recent advances in parameter-efficient fine-tuning (PEFT) techniques have demonstrated the feasibility of learning representations while introducing a limited number of new parameters to adapt a frozen pre-trained model to downstream tasks. Its core idea is to redesign downstream tasks to align closely with those addressed during pre-training, ensuring that large-scale pre-trained knowledge is effectively retained. Owing to PEFT, large-scale pre-trained visual-language models (e.g., CLIP (Radford et al., 2021)) have achieved impressive results in various visual tasks. However, existing PEFT methods (Zhou et al., 2022a, b) are designed for multimodal models to capture high-level category semantics rather than subtle subcategory differences, so directly applying existing PEFT techniques to FGIR without task-specific adaptation tends to be suboptimal, due to their limited focus on capturing subtle intra-class variations. Thus, developing an efficient fine-tuning strategy tailored for fine-grained visual models is essential to avoid suboptimal convergence when fine-tuning the entire FGIR model.
@@
To this end, we design a novel adaptation framework for FGIR, which efficiently unleashes the power of pre-trained model, as shown in Fig. 1(b). Technically, we propose Dual-Vision Adaptation (DVA) to resolve the conflict between preserving pre-trained knowledge and acquiring fine-grained discriminative power. By maintaining the frozen backbone, DVA establishes category-sensitive perception through sample-feature co-adaptation, inherently avoiding the optimization dilemma of conventional full fine-tuning. Specifically, the Object-Perceptual Adaption (OPA) is a proposed sample-adapted strategy for the visual foundation model, encouraging the pre-trained model to capture critical object regions and locate subclass-specific differences. Meanwhile, as a feature adaptation process, the proposed In-Context Adaption (ICA) aims to focus on fine-grained features that contribute to category prediction through lightweight learnable parameters. In this way, ICA can adapt the feature content in a direction that is conducive to category recognition, which makes the FGIR task with such adjusted features close to the task solved during the original pre-training. Nevertheless, a non-negligible problem is that the integration between the visual foundation model and the frozen pre-trained backbone introduces prohibitive computational overhead during inference. To address this critical efficiency bottleneck while retaining fine-grained discriminative power, we devise a Discriminative Perceptual Transfer (DPT) module that transfers the discriminative knowledge from OPA to the backbone via a distillation mechanism.
@@
  • With only 3.5% of the tunable parameters compared to full fine-tuning, DVA achieves competitive performance on three in-distribution and three out-of-distribution fine-grained datasets.
2. Related Works
2.1. Fine-Grained Image Retrieval
@@
Existing methods for fine-grained image retrieval (FGIR) can be categorized as encoding-based or localization-based. The encoding-based methods (Movshovitz-Attias et al., 2017; Teh et al., 2020; Lim et al., 2022; Ermolov et al., 2022; Jiang et al., 2024c, a) aim to learn an embedding space in which samples of a similar subcategory are attracted and samples of different subcategories are repelled. The methods can be decomposed into roughly two components: the image encoder maps images into an embedding space, and the metric method ensures that samples from the same subcategories are grouped closely, while samples from different subcategories are separated. While these studies have achieved significant achievements, they primarily concentrate on optimizing image-level features that include numerous noisy and non-discriminatory details. Therefore, the localization-based methods (Han et al., 2018; Tang et al., 2020; Wang et al., 2023b; Wei et al., 2017; Moskvyak et al., 2021; Wang et al., 2022) are proposed, which focus on training a subnetwork for locating discriminative regions or devising an effective strategy for extracting attractive object features to facilitate the retrieval task. Unlike these approaches, our method considers the specific characteristics of the FGIR tasks, offering guidance for designing high-performance retrieval models.
2.2. Parameter-Efficient Fine Tuning
@@
Parameter-efficient fine-tuning (Hu et al., 2022) in NLP reformulates downstream tasks as language modeling problems, allowing pre-trained language models to adapt more efficiently to new tasks. As a result, these techniques are now widely used in various NLP applications, including language understanding and generation. Recently, parameter-efficient fine-tuning has also been applied to multimodal computer vision (Zhou et al., 2022b, a; Jiang et al., 2024b; Cao et al., 2023, 2024b, 2024a). However, existing methods mainly extend the capabilities of language-based models, making them unsuitable for direct use with pre-trained vision models. Additionally, some recent approaches (Chen et al., 2022a; Gao et al., 2024) introduce a small number of learnable parameters to guide pre-trained models in general vision tasks. However, their fine-tuning strategies focus on capturing category-level semantics rather than the fine-grained visual differences needed to distinguish similar objects. To address this limitation, we propose a parameter-efficient DVA that incorporates feature and sample adaptation, enabling the frozen pre-trained vision model to perform FGIR tasks effectively.
2.3. Knowledge Distillation
Knowledge Distillation (KD) is dedicated to compressing the informative knowledge from a large and computationally expensive model (i.e, teacher model) to a small and computationally efficient model (i.e, student model) (Hinton et al., 2015; Fang et al., 2024). Given the strong knowledge transfer ability, it is widely applied to natural language processing, computer vision and other fields (Wang and Yoon, 2021; Hu et al., 2023). Most classification-based KD approaches explore enhancing the student model by mimicking the teacher model’s predictions or output distribution. Additionally, researchers have delved into exploring KD for image retrieval by leveraging distances between samples, such as learning to rank (Chen et al., 2018) and regression on quantities involving one or more pairs, like distances (Wang et al., 2023a) or angles (Park et al., 2019). In this paper, we transfer discriminative knowledge that aids fine-grained image retrieval via proxy features.
3. Methods
@@
Figure 2. The overall framework of our proposed DVA. It consists of three crucial components, $i.e$, Object-Perceptual Adaption (OPA), In-Context Adaption (ICA) and Discriminative Perceptual Transfer (DPT). OPA enhances the image encoder’s ability to capture discriminative object regions and ICA dynamically adapts fine-grained features while suppressing irrelevant background retained in the frozen representations. Finally, DPT inject a discriminative awareness into the image encoder via distillation, enabling auxiliary-free inference while preserving pre-trained knowledge.
3.1. Image Encoder
For the input image $\textbf{I}\in\mathbb{R}^{3\times H\times W}$, the image encoder will first split it into $N=N_{h}\times N_{w}$ non-overlapping patches of size $P\times P$. Here, $N_{h}=\frac{H}{P}$ and $N_{w}=\frac{W}{P}$. Subsequently, these patches are transformed into embedding tokens $\textbf{E}=[\textbf{E}^{1},\textbf{E}^{2},\dots,\textbf{E}^{N}]\in\mathbb{R}^{N\times D}$ using learnable linear projection, where $D$ denotes the dimension of each token. Finally, the embedding tokens E are concatenated with a class token $\textbf{E}^{\mathrm{{cls}}}\in\mathbb{R}^{D}$, and the position information of the tokens is retained through combined a learnable position embedding $\textbf{E}_{\mathrm{pos}}$ to form the initial input token sequence as $\textbf{E}_{0}=[\textbf{E}^{\mathrm{cls}},\textbf{E}^{1},\textbf{E}^{2},\dots,\textbf{E}^{N}]+\textbf{E}_{\mathrm{pos}}$. The image encoder comprises $L$ layers, each incorporating multi-head self-attention (MHSA) and a multi-layer perception (MLP) block (Dosovitskiy et al., 2020). Therefore, considering the input $\textbf{E}_{i-1}$ of the $i$-th layer, the output $\textbf{E}_{i}$ can be obtained as follows:
(1) $\mathbf{E}_{i}^{\prime}$ $=\textrm{MHSA}(\textrm{LN}(\mathbf{E}_{i-1}))+\mathbf{E}_{i-1},$
(2) $\mathbf{E}_{i}$ $=\textrm{MLP}(\textrm{LN}(\mathbf{E}_{i}^{\prime}))+\mathbf{E}_{i}^{\prime},$
where $i=\{1,2,\dots,L\}$, and $\textrm{LN}(\cdot)$ denotes the layer normalization operation. More specifically, the MHSA module projects $\mathbf{E}_{i-1}$ into queries $\mathbf{Q}_{i-1}$, keys $\mathbf{K}_{i-1}$, and values $\mathbf{V}_{i-1}$, each $\in\mathbb{R}^{(N+1)\times D}$, with a attention projector.
(3) $\mathbf{Q}_{i-1},\mathbf{K}_{i-1},\mathbf{V}_{i-1}=\textrm{AttentionProjector}(\textrm{LN}(\mathbf{E}_{i-1})).$
These projections are then split into $H$ heads to achieve parallel attention, followed by an output projector to obtain the output $\mathbf{E}^{mhsa}_{i-1}$ of MHSA:
(4) $\mathbf{E}^{mhsa}_{i-1}=\textrm{OutputProjector}(\mathrm{softmax}\Bigl{(}\frac{\mathbf{Q}_{i-1}\mathbf{K}_{i-1}^{\top}}{\sqrt{D_{h}}}\Bigr{)}\mathbf{V}_{i-1}),$
where $D_{h}=\frac{D}{H}$.
Finally, we serve the class token $\textbf{E}^{\mathrm{cls}}$ from $\textbf{E}_{L}$ at the $L$-th layer as the retrieval embedding $\textbf{E}^{R}$.
3.2. Dual-Vision Adaptation
3.2.1. Object-Perceptual Adaption
@@
Subtle yet discriminative discrepancies are widely recognized to be significant for fine-grained understanding (Fang et al., 2024; Wang et al., 2023b; Chen et al., 2022b). However, the vanilla ViT model is originally designed to identify different species (e.g., cats, dogs, and birds), rather than to exploit subtle differences between subcategories within a species. Therefore, without task-specific adaptation, the vanilla ViT has limited focus on capturing subtle intra-class variations. To alleviate this situation, we design an Object-Perceptual Adaption (OPA) that only modifies the visual input of the image encoder (ViT) to assist the model in perceiving discriminative differences in fine-grained images. Specifically, OPA consists of object perception and discriminative perception. The former better perceives objects through the background separated by the visual foundation model, and the latter perceives category differences through the critical foreground separated by the visual foundation model.
@@
Discriminative Perception. For fine-grained understanding, there have been many works (Wang et al., 2023b; Jiang et al., 2024c) demonstrating that the background can perturb the model’s perception of discriminative regions. Therefore, in contrast to object perception, discriminative perception focuses on objects in the approach image, thereby helping to explore discriminative features through critical regions. We reuse the object perception process, the difference is that we remove the background area and keep the core object foreground, obtaining the discriminative image $\mathbf{I}_{\mathrm{d}}$. To ensure objects are fully contained in the image and free from deformation, we follow DVF (Jiang et al., 2024c) to filter out low-confidence detection results, and finally apply post-processing to preserve the original shapes of the object.
3.2.2. In-Context Adaptation
@@
Fine-Tuning Preferences Analysis of ViT: Conventional methods (Ermolov et al., 2022; Fang et al., 2024) usually require fine-tuning the entire network, but this approach ignores the large-scale pre-trained knowledge present in the pre-trained parameters. We hypothesize that fine-grained image retrieval on downstream datasets can be learned and understood by specific modules within the network. To verify this hypothesis, we conducted an in-depth analysis of the fine-tune preferences of ViT.
@@
We first divided the trainable parameters of transformer layers into three parts: the attention projector, output projector and MLP. Subsequently, we separately fine-tune these three parts and ViT as a whole, and use the fine-tuned models to test the results on the CUB-200-2011 (Branson et al., 2014) dataset. As shown in Table 1, we observed that only by fine-tuning part modules in ViT, we can achieve performance similar to that of fine-tuning the entire model. Meanwhile, fine-tuning the attention projector yields the best results. This is intuitive since the self-attention mechanism is central to ViT, and the attention projector directly determines output. This experiment further validates our hypothesis that the image retrieval task can be learned by a specific network module, namely attention projector.
@@
Based on this insight, we attempt to tune only the attention projector of the ViT. However, this approach still destroys the knowledge in the pre-trained parameters and still requires learning a large number of parameters (22.5M). To address this issue, we keep the attention projector frozen and propose a lighter-weight learnable in-context adaptation module in parallel with the attention projector to learn the specific knowledge required for fine-grained image retrieval. Specifically, the in-context adaptation module is designed to be a bottleneck structure for limiting the number of parameters purpose, which includes a down-projection layer with parameters $\mathbf{W}_{down}\in\mathbb{R}^{D\times d}$, an up-projection layer with parameters $\mathbf{W}_{up}\in\mathbb{R}^{d\times D}$, in this work, we set $d=16$. For a specific input feature $\mathbf{x}$, the in-context adaptation module produces the in-context features, $\mathbf{x}_{ic}$, formally via:
(5) $\mathbf{x}_{ic}=\textrm{LN}(\mathbf{x})\cdot\mathbf{W}_{down}\cdot\mathbf{W}_{up}.$
@@
Then, we can replace the original attention map output with a feature that contains both pre-trained knowledge and dataset-specific knowledge through residual connections:
(6) $\mathbf{\hat{Q}},\mathbf{\hat{K}},\mathbf{\hat{V}}=\textrm{AttentionProjector}(\textrm{LN}(\mathbf{E}))+\textrm{IC}(\mathbf{E}),$
@@
where $\textrm{IC}(\cdot)$ represents the multiple in-context adaptation modules of $\mathbf{E}$ corresponding to the attention projectors that generate $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$. In this paper, we only assign in-context adaptation modules to the attention projectors that generate $\mathbf{Q}$ and $\mathbf{K}$. For more experimental analysis, please see Section 4.4.3.
Table 1. Fine-Tuning Preferences Analysis of ViT on the CUB-200-2011 dataset.
Methods R@1 R@2 R@4 R@8 Prams (M)
Fine-tuning 82.7 88.7 92.5 95.2 85.6
Attention Projector 82.7 89.0 92.8 95.2 21.5
Output Projector 82.6 88.8 92.6 95.0 7.3
MLP 82.0 88.9 92.9 95.2 56.8
3.3. Discriminative Perceptual Transfer
@@
With the cooperation of discriminative perception, the fine-grained understanding ability of the ViT model is enhanced. However, the extra visual foundation model is time-consuming and memory-demanding for retrieval evaluation. In the classification field (Fang et al., 2024), network distillation has been shown to be one of the solutions to this problem. Inspired by network distillation, we propose a discriminative perceptual transfer to extend the knowledge distillation theory to retrieval tasks. Specifically, we learn discriminative category proxies as carriers of discriminative knowledge and transfer their discriminative knowledge using the same learning strategy.
@@
Category Proxy Learning. We encode both the discrimination and distribution of labelled instances via proxy-guided learning (Kim et al., 2020). Then, for the retrieval embedding $\textbf{E}^{R}$ , we compute the proxy learning loss as follow:

### v4

@@ Abstract
Fine-Grained Image Retrieval (FGIR) faces challenges in learning discriminative visual representations to retrieve images with similar fine-grained features. Current leading FGIR solutions typically follow two regimes: enforce pairwise similarity constraints in the semantic embedding space, or incorporate a localization sub-network to fine-tune the entire model. However, such two regimes tend to overfit the training data while forgetting the knowledge gained from large-scale pre-training, thus reducing their generalization ability. In this paper, we propose a Dual-Vision Adaptation (DVA) approach for FGIR, which guides the frozen pre-trained model to perform FGIR through collaborative sample and feature adaptation. Specifically, we design Object-Perceptual Adaptation, which modifies input samples to help the pre-trained model perceive critical objects and elements within objects that are helpful for category prediction. Meanwhile, we propose In-Context Adaptation, which introduces a small set of parameters for feature adaptation without modifying the pre-trained parameters. This makes the FGIR task using these adapted features closer to the task solved during the pre-training. Additionally, to balance retrieval efficiency and performance, we propose Discrimination Perception Transfer to transfer the discriminative knowledge in the object-perceptual adaptation to the image encoder using the knowledge distillation mechanism. Extensive experiments show that DVA performs well on three fine-grained datasets.
Introduction
@@
Figure 1: (a) previous fine-tuning methods. (b) our dual-vision adaptation method. Our approach designs the collaborative sample and feature adaptation to exploit category-specific differences. This dual strategy enables model to sustain broad representation capabilities from pre-training data while dynamically adjusting its adaptability to fine-grained data.
Unlike general image retrieval, fine-grained image retrieval (FGIR) attempts to retrieve images belonging to the same subcategory as the query image from a database within a broader meta category (i.e., birds, cars) (WeiLWZ17). It has been extensively applied across various domains, including intelligent transportation (ramdurai2025large) and biodiversity monitoring (vendrow2024inquire). However, FGIR presents significant challenges due to two inherent difficulties: i) subtle visual distinctions between different categories, and ii) significant appearance variations within the same category. This task usually requires models to simultaneously localize discriminative regions and identify minute visual differences, creating a dual optimization dilemma. Therefore, the key to FGIR is learning discriminative and generalizable embeddings to identify visually similar objects accurately.
Conventional FGIR methods typically adopt two primary strategies: encoding-based (pnca; pnca++; Hyp-ViT; hist) and localization-based (ZhengJSWHY18; frpt; WeiLWZ17; wang2022coarse). Encoding-based methods primarily learn image-level features but struggle to suppress background and irrelevant information interference. In contrast, localization-based methods enhance feature extraction within the encoder to capture subtle differences among subcategories, often by focusing on distinctive object regions. Although both strategies capture discriminative embedding learning for fine-grained retrieval, their dependence on full-parameter fine-tuning introduces prohibitive computational costs and limits cross-task generalization, as shown in Fig. 1(a). Simultaneously, fine-tuning the entire model destroys the knowledge in the large-scale pre-trained parameters and may lead to suboptimal convergence, which ultimately limits retrieval performance (frpt). This raises the question: can discriminative representations be learned for FGIR tasks without fine-tuning the model’s pre-trained parameters?
@@
Recent advancements in parameter-efficient fine-tuning (PEFT) techniques have demonstrated the feasibility of learning representations while introducing a limited number of new parameters to adapt a frozen pre-trained model to downstream tasks. Its core idea is to redesign downstream tasks to align closely with those addressed during pre-training, ensuring that large-scale pre-trained knowledge is effectively retained. Owing to PEFT, large-scale pre-trained visual-language models (e.g., CLIP (clip)) have achieved impressive results in various visual tasks. However, existing PEFT methods (zhou2022cocoop; zhou2022coop) are designed for multimodal models to capture high-level category semantics rather than subtle subcategory differences, so directly applying existing PEFT techniques to FGIR without task-specific adaptation tends to be suboptimal, due to their limited focus on capturing subtle intra-class variations. Therefore, it is crucial to develop an efficient fine-tuning strategy specifically designed for fine-grained visual models, as this can prevent suboptimal convergence that may occur when fine-tuning the entire FGIR model.
@@
To this end, we design a novel adaptation framework for FGIR, which efficiently unleashes the power of pre-trained model, as shown in Fig. 1(b). Technically, we propose Dual-Vision Adaptation (DVA) to resolve the conflict between preserving pre-trained knowledge and acquiring fine-grained discriminative power. By maintaining the frozen backbone, DVA establishes category-sensitive perception through sample-feature co-adaptation, inherently avoiding the optimization dilemma of conventional full fine-tuning. Specifically, the Object-Perceptual Adaption (OPA) is a proposed sample-adapted strategy for the visual foundation model, encouraging the pre-trained model to capture critical object regions and locate subclass-specific differences. Meanwhile, as a feature adaptation process, the proposed In-Context Adaption (ICA) aims to focus on fine-grained features that contribute to category prediction through lightweight learnable parameters. In this way, ICA can adapt the feature content in a direction that is conducive to category recognition, which makes the FGIR task with such adapted features close to the task solved during the original pre-training. Nevertheless, a non-negligible problem is that the integration between the visual foundation model and the frozen pre-trained backbone introduces prohibitive computational overhead during inference. To address this critical efficiency bottleneck while retaining fine-grained discriminative power, we devise a Discriminative Perceptual Transfer (DPT) module that transfers the discriminative knowledge from OPA to the backbone via a distillation mechanism.
@@
  • With only 0.68% of the tunable parameters compared to full fine-tuning, DVA achieves competitive performance on three fine-grained datasets.
Related Works
Fine-Grained Image Retrieval
@@
Existing methods for fine-grained image retrieval (FGIR) can be categorized as encoding-based or localization-based. The encoding-based methods (pnca; pnca++; hist; Hyp-ViT; jiang2024dvf; jiang2024global; tang2020blockmix) aim to learn an embedding space in which samples of a similar subcategory are attracted and samples of different subcategories are repelled. The methods can be decomposed into roughly two components: the image encoder maps images into an embedding space, and the metric method ensures that samples from the same subcategories are grouped closely, while samples from different subcategories are separated. While these studies have achieved significant achievements, they primarily concentrate on optimizing image-level features that include numerous noisy and non-discriminatory details. Therefore, the localization-based methods (HanGZZ18; TangLPT20; frpt; WeiLWZ17; moskvyak2021keypoint; tang2022learning) are proposed, which focus on training a subnetwork for locating discriminative regions or devising an effective strategy for extracting attractive object features to facilitate the retrieval task. Unlike these approaches, our method considers the specific characteristics of the FGIR tasks, offering guidance for designing high-performance retrieval models.
Parameter-Efficient Fine Tuning
@@
Parameter-efficient fine-tuning (hu2022lora) in NLP reformulates downstream tasks as language modeling problems, allowing pre-trained language models to adapt more efficiently to new tasks. As a result, these techniques are now widely used in various NLP applications, including language understanding and generation. Recently, parameter-efficient fine-tuning has also been applied to multimodal computer vision (zhou2022coop; xing2025inv; mpfgvc; tang2025connecting; cao2024eventcrab; gao2025styleshot). However, existing methods primarily enhance language-based models, making them unsuitable for direct application to pre-trained vision models. Additionally, some recent approaches (chen2022adaptformer; gao2025faceshot) introduce a small number of learnable parameters to guide pre-trained models in general vision tasks. However, their fine-tuning strategies focus on capturing category-level semantics rather than the fine-grained visual differences needed to distinguish similar objects. To address this limitation, we propose a parameter-efficient DVA that incorporates sample and feature adaptation, enabling the frozen pre-trained vision model to perform FGIR tasks effectively.
@@
Figure 2: DVA consists of three essential modules: the Object-Perceptual Adaptation module to enhance the encoder’s ability to focus on discriminative object regions, the In-Context Adaptation module to dynamically refine fine-grained features while suppressing irrelevant background retained in frozen representations, and the Discriminative Perceptual Transfer module to distill discriminative awareness into the encoder, enabling auxiliary-free inference while preserving pre-trained knowledge.
Knowledge Distillation
Knowledge Distillation (KD) is dedicated to compressing the informative knowledge from a large and computationally expensive model (i.e, teacher model) to a small and computationally efficient model (i.e, student model) (hinton2015distilling; csdnet; cao2025exploiting). Given the strong knowledge transfer ability, it is widely applied to natural language processing, computer vision and other fields (wang2021knowledge; tang2023m3net; tang2024divide). Most classification-based KD approaches explore enhancing the student model by mimicking the teacher model’s predictions or output distribution. Additionally, researchers have delved into exploring KD for image retrieval by leveraging distances between samples, such as learning to rank (chen2018darkrank) and regression on quantities involving one or more pairs, like distances (pleor) or angles (park2019relational). In this paper, we transfer discriminative knowledge that aids fine-grained image retrieval via proxy features.
Method
Dual-Vision Adaptation
Object-Perceptual Adaption.
@@
Subtle yet discriminative discrepancies are widely recognized to be significant for fine-grained understanding (csdnet; frpt; cao2024eventcrab). However, the vanilla ViT model is originally designed to identify different species (e.g., cats, dogs, and birds), rather than to exploit subtle differences between subcategories within a species. Therefore, without task-specific adaptation, the vanilla ViT has limited focus on capturing subtle intra-class variations. To alleviate this situation, we design an Object-Perceptual Adaption (OPA) that only modifies the visual input of the image encoder (ViT) to assist the model in perceiving discriminative differences in fine-grained images. Specifically, OPA consists of two components: discriminative perception and object perception. The former enhances category discrimination by focusing on foreground regions identified by the visual foundation model, while the latter improves object perception by utilizing background information separated by the same model.
Discriminative Perception. For fine-grained understanding, there have been many works (xing2024csgo; jiang2024dvf) demonstrating that the background can perturb the model’s perception of discriminative regions. Therefore, we design discriminative perception to help the model focus on objects in the image, thereby exploring discriminative features through critical regions. We use visual foundation model to extract object regions from training images, keep the core object foreground, obtaining the discriminative image $\mathbf{I}_{\mathrm{d}}$. Specifically, each training image along with the super-class name (e.g., bird for CUB-200-2011 (cub)) of the image is fed to open-vocabulary detector GroundingDINO (grounding-dino) to obtain the bounding box that contains the foreground object of the category. Therefore, there is no risk of information leakage for FGIR. Then, the region containing the foreground object is extracted from the image as the discriminative image $\mathbf{I}_{\mathrm{d}}$. To ensure objects are fully contained in the image and free from deformation, we filter out low-confidence detection results, and finally apply padding on the short sides of the image to preserve the original shape of the object.
@@
In-Context Adaptation.
@@
Conventional methods (Hyp-ViT; csdnet; cao2024adafpp) usually require fine-tuning the entire network, but this approach ignores the large-scale pre-trained knowledge present in the pre-trained parameters. We hypothesize that FGIR on downstream datasets can be learned and understood by specific modules within the network. To verify this hypothesis, we conducte an in-depth analysis of the fine-tune preferences of ViT.
@@
We first divide the trainable parameters of transformer layers into three parts: the attention projector, output projector and MLP. Subsequently, we separately fine-tune these three parts and ViT as a whole and evaluate on the CUB-200-2011 (cub) dataset. As shown in Table 1, we found that by fine-tuning only specific modules within ViT, it is possible to achieve performance comparable to that of fine-tuning the entire model. Meanwhile, fine-tuning the attention projector yields the best performance. This is intuitive, as the self-attention mechanism is central to ViT, and the attention projector enhances its ability to encode contextual information. This experiment further validates our hypothesis that the FGIR task can be learned by a specific network module, namely attention projector.
@@
Based on this insight, we attempt to design a in-context adaptation module with tune only the attention projector of the ViT. However, this approach still destroys the knowledge in the pre-trained parameters and still requires learning a large number of parameters (22.5M). To address this issue, we keep the attention projector frozen and propose a lighter-weight learnable in-context adaptation module in parallel with the attention projector to learn the specific knowledge required for FGIR. Specifically, the in-context adaptation module is designed to be a bottleneck structure for limiting the number of parameters purpose, which includes a down-projection layer with parameters $\mathbf{W}_{down}\in\mathbb{R}^{D\times d}$, an up-projection layer with parameters $\mathbf{W}_{up}\in\mathbb{R}^{d\times D}$, in this work, we set $d=16$. For a specific input feature $\mathbf{x}$, the in-context adaptation module produces the in-context features, $\mathbf{x}_{ic}$, formally via:
$\mathbf{x}_{ic}=\textrm{LN}(\mathbf{x})\cdot\mathbf{W}_{down}\cdot\mathbf{W}_{up}.$ (1)
@@
Then, we can replace the original attention projector output with a feature that contains both pre-trained knowledge and task-specific knowledge through residual connections:
$\mathbf{\hat{Q}}=\textrm{AttentionProjector}_{q}(\textrm{LN}(\mathbf{E}))+\textrm{IC}_{q}(\mathbf{E}),$ (2)
@@
where $\textrm{IC}_{q}(\cdot)$ represents the in-context adaptation module of $\mathbf{E}$ corresponding to the $\textrm{AttentionProjector}_{q}(\cdot)$ that generate query $\mathbf{Q}$. In this paper, we only assign in-context adaptation modules to the attention projectors that generate $\mathbf{Q}$ and $\mathbf{K}$. For experimental analysis, please see Experimental section.
Methods R@1 R@2 R@4 Prams (M)
Non-fine-tuned 72.5 82.6 89.1 0.0
Fine-tuning 82.7 88.7 92.5 85.6
Attention Projector 82.7 89.0 92.8 21.5
Output Projector 82.6 88.8 92.6 7.3
MLP 82.0 88.9 92.9 56.8
Table 1: Fine-Tuning Preferences Analysis of ViT on the CUB-200-2011 dataset.
Discriminative Perceptual Transfer
@@
With the cooperation of discriminative perception, the fine-grained understanding ability of the ViT model is enhanced. However, the extra visual foundation model is time-consuming and memory-demanding for retrieval evaluation. In the classification field (hinton2015distilling), network distillation has been shown to be one of the solutions to this problem. Inspired by network distillation, we propose a discriminative perceptual transfer to extend the knowledge distillation theory to retrieval tasks to achieve an efficiency-performance tradeoff. Specifically, we separate discriminative perception into a separate branch, exploit a category proxy learning strategy to learn discriminative category proxy as carriers of discriminative knowledge, and transfer their discriminative knowledge.
@@
We encode both the discrimination and distribution of labelled instances via proxy-guided learning (proxy). Then, for the retrieval embedding $\textbf{E}^{R}$ , we compute the proxy learning loss as follow:
$\mathcal{L}_{proxy}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\textbf{E}^{R}\|,\|\textbf{c}\|))}\right),$ (3)
where $d(\|\textbf{E}^{R}\|,\|\textbf{c}\|)$ represents the distance between $\|\textbf{E}^{R}\|$ and $\|\textbf{c}\|$, c denotes the category proxy corresponding to retrieval embedding $\textbf{E}^{R}$, C is the category proxy set of the discriminative perception images $\mathbf{I}_{d}$, and $\|\cdot\|$ denotes the $L^{2}$-Norm.
CUB-200-2011 Stanford Cars
Method R@1 R@2 R@4 R@8 R@1 R@2 R@4 R@8
Proxy-Anchor (proxy) 80.4 85.7 89.3 92.3 77.2 83.0 87.2 90.2
HIST (hist) 75.6 83.0 88.3 91.9 89.2 93.4 95.9 97.6
PNCA++  (pnca++) 80.4 85.7 89.3 92.3 86.4 92.3 96.0 97.8
Hyp-ViT (Hyp-ViT) 84.2 91.0 94.3 96.0 76.7 85.2 90.8 94.7
DVA (Ours) 85.2 91.4 94.6 96.2 88.0 93.2 97.0 98.5
Table 2: Comparison with state-of-the-art methods in the closed-set setting on CUB-200-2011 and Stanford Cars. The best result is shown in bold, and the second-best result is underlined.
With the discriminative category proxy set C, and the retrieval embeddings $\hat{\textbf{E}}^{R}$ of the origin training images are used for distillation:
$\mathcal{L}_{dpt}=-\textrm{log}\left(\frac{\textrm{exp}(-d(\|\hat{\textbf{E}}^{R}\|,\|\textbf{c}\|))}{\sum_{\textbf{c}\in\textbf{C}}\textrm{exp}(-d(\|\hat{\textbf{E}}^{R}\|,\|\textbf{c}\|))}\right).$ (4)
Overall Function
As illustrated in Fig. 2, our objective function comprises two components: the proxy learning loss $\hat{\mathcal{L}}_{proxy}$ and the distillation loss $\mathcal{L}_{dpt}$. Finally, the total loss for our framework can be defined as: